MULTIPLYING FRACTIONS and MIXED NUMBERS
1/3 + 3/5y = 2/5

Im uber stumped.

Answer by stanbon(75887)

1/3 + (3/5)y = 2/5
---------
(3/5)y = (2/5)-(1/3)
-----
(3/5)y = (6-5)/15
------
(3/5)y = 1/15
y = (5/3)(1/15)
-----
y = 1
============
Cheers,
Stan H.